# Open-World Skill Discovery from Unsegmented Demonstrations

### v1

@@ Abstract
Learning skills in open-world environments is essential for developing agents capable of handling a variety of tasks by combining basic skills. Online demonstration videos are typically long but unsegmented, making them difficult to segment and label with skill identifiers. Unlike existing methods that rely on sequence sampling or human labeling, we have developed a self-supervised learning-based approach to segment these long videos into a series of semantic-aware and skill-consistent segments. Drawing inspiration from human cognitive event segmentation theory, we introduce Skill Boundary Detection (SBD), an annotation-free temporal video segmentation algorithm. SBD detects skill boundaries in a video by leveraging prediction errors from a pretrained unconditional action-prediction model. This approach is based on the assumption that a significant increase in prediction error indicates a shift in the skill being executed. We evaluated our method in Minecraft, a rich open-world simulator with extensive gameplay videos available online. Our SBD-generated segments improved the average performance of conditioned policies by 63.7% and 52.1% on short-term atomic skill tasks, and their corresponding hierarchical agents by 11.3% and 20.8% on long-horizon tasks. Our method can leverage the diverse YouTube videos to train instruction-following agents. The project page can be found in https://craftjarvis.github.io/SkillDiscovery/.
[…]
Yitao Liang <yitaol@pku.edu.cn>
[…]
Jingwen Deng <dengjingwen@stu.pku.edu.cn>, Zihao Wang <zhwang@stu.pku.edu.cn>, Anji Liu <liuanji@cs.ucla.edu>
@@ 1 Introduction
Sequential sampling [Uncaptioned image] rule-based
@@
Table 1: Comparisons between existing segmentation methods and our method SBD. Existing methods usually rely on human-designed rules, while our method is learning-based. Sequential sampling can result in a single skill spanning different segments or multiple skills located within one segment. Reward-driven methods require additional reward information, which is challenging for human annotators to label. Top-down methods often result in limited skill diversity and high computation costs. Bottom-up methods are limited in fully observable environments and hard in visually partially observable environments. [Uncaptioned image] are visual observations and [Uncaptioned image] are actions.
Most existing LLM-based instruction-following agents adopt a two-layer structure of planner and controller to complete tasks, which first convert the instruction into atom skills through the planner, and then use a conditioned policy to convert it into actions based on the current observation to interact with the environment (Wang et al., 2025, 2023b, 2024b; Li et al., 2024; Wang et al., 2024a; Cheng et al., 2024). When training these agents, we need to first convert long sequences into a series of short atom skill sequences, and then train the controller and planner separately. However, interactions in real-world videos are often long-term and unsegmented with detailed language skills.
@@
Learning skills from long-sequence videos is critical for building such hierarchical agents. However, the concept of a “skill” is ill-defined and varies widely across domains such as video gaming (Cai et al., 2023b; Wang et al., 2024a; Hafner et al., 2023; Zhang et al., 2024), robotic control (Zitkovich et al., 2023), and autonomous driving (Wang et al., 2023a). This ambiguity, along with the immense diversity of skills in open worlds, makes skill learning particularly challenging, especially in partially observable settings. To enable open-world skill learning from unsegmented demonstrations (continuous streams of observation-action pairs without explicit labels), the first challenge is to segment these streams into semantically meaningful, self-contained skills.
@@
We list the existing segmentation methods in Table 1, including the plain sequential sampling, reward-driven methods, top-down, and bottom-up methods. The native sequential sampling segmentation methods (Cai et al., 2023b; Lifshitz et al., 2023), divide videos into segments of predefined lengths (e.g., fixed length, uniform distribution). However, these methods do not ensure that each segment contains a distinct skill. Additionally, predefined lengths may not match the actual distribution of skill lengths in real-world scenarios (see Section 4.4). Reward-driven methods (Sutton et al., 1999) discover skills through the environment’s reward signal. It is limited by its inability to capture skills that lack associated rewards and by the risk of splitting a single skill into multiple segments when rewards are repeatedly gained during execution. Top-down methods (Shiarlis et al., 2018; Belkhale et al., 2024) rely on predefined skill sets from human experts. They use manual labeling or supervised learning to segment videos. Although this approach can produce reasonable results, it is expensive and limited by the narrow range of predefined skills. Bottom-up methods (Zhu et al., 2022; Pertsch et al., 2025) use algorithms such as agglomerative clustering or byte-pair encoding (BPE) (Gage, 1994) to split action sequences. However, they struggle in partially observable settings where both observations and actions must be considered because these algorithms cannot deal with observations due to the high dimensionality. All the above methods usually rely on human-designed rules to segment the unsegmented videos, which highlights the need for an effective and label-free method that can automatically segment skills from unsegmented demonstrations in open worlds.
@@
Inspired by event segmentation theories (EST) (Zacks et al., 2007), which propose that humans naturally partition continuous experiences into discrete events when prediction errors in perceptual expectations rise, we introduce Skill Boundary Detection (SBD)—a method for autonomously identifying potential skill boundaries within extended trajectories. SBD employs a predictive model trained on a dataset of unsegmented videos to forecast future actions based on past observations, effectively capturing temporal dependencies (Baker et al., 2022). We then use this pretrained model to make predictions on an unsegmented video and compare them to the ground truth actions. A significant increase in prediction error indicates a shift in the skill being executed (Theorem 3.4), enabling us to detect boundaries between different skills in the video. SBD relies on self-supervised learning, removing the need for additional human labeling. This allows SBD to utilize a wide range of YouTube videos to train instruction-following agents.
@@
We evaluate SBD in open-world Minecraft (Guss et al., 2019; Fan et al., 2022; Lin et al., 2023). First, we apply SBD to create a segmented Minecraft video dataset. We then train the video-conditioned policy (Cai et al., 2023b) and the language-conditioned policy (Lifshitz et al., 2023) on this dataset. We evaluate these policies on diverse Minecraft skills benchmark (Lin et al., 2023). The results show significant improvements compared to the existing policies (Cai et al., 2023b; Lifshitz et al., 2023), with performance increases of 63.7% and 52.1% compared to the original methods, respectively. We also test their corresponding hierarchical agents using behavior cloning and in-context learning, achieving performance increases of 11.3% and 20.8%, respectively. These findings underscore the effectiveness of our proposed method for skill discovery from unsegmented demonstrations and its potential to advance open-world skill learning.
@@
Figure 1: Pipeline of our method SBD for discovering skills from unsegmented demonstration videos. Stage I: An unconditional Transformer-XL based policy model (Dai et al., 2019; Baker et al., 2022) is pretrained on an unsegmented dataset to predict future actions (labeled by an inverse dynamics model) based on past observations. Stage II: The pretrained unconditional policy will produce a high predicted action loss when encountering uncertain observations (e.g., deciding whether to kill a new sheep) in open worlds. These timesteps should be marked as skill boundaries, indicating the need for additional instructions to control behaviors. We segment the long unsegmented videos into a series of short atomic skill demonstrations. Stage III: We train a conditional Transformer-XL based policy model on the segmented dataset to master a variety of atomic skills. Stage IV: Finally, we use hierarchical methods (a combination of vision-language models and the conditional policy) to model the long demonstration and follow long-horizon instructions.
@@ 2 Problem Formulation
Behavioral cloning learns a policy $\pi$ that maps observations $o$ from the observation space $\mathcal{O}$ to actions $a$ in the action space $\mathcal{A}$ using supervised learning on observation-action pairs $(o_{i},a_{i})_{i=1}^{T}$ obtained from expert demonstrations. In partially observable environments, past observations $o_{1:t}$ are often used instead of the current observation $o_{t}$ as input to the policy $\pi$. Additionally, to allow the agent to dynamically adjust its behavior based on its objective, recent advances (Khandelwal et al., 2021; Du et al., 2021; Xie et al., 2023) focus on learning goal-conditioned policies $\pi(o_{1:t},g)$, where $g$ represents the goal. This type of goal-conditioned policy can be considered a “skill”.
@@
Directly learning policy $\pi(o_{1:t},g)$ is challenging in complex open-world environments like Minecraft. Goals such as build a house or mine diamond blocks involve numerous intermediate steps, making it computationally and conceptually difficult to model an agent’s behavior without breaking the goal into manageable parts. To tackle this, hierarchical agent architectures are often used (Wang et al., 2024a; Driess et al., 2023; Cai et al., 2024a; Iqbal et al., 2022). These architectures consist of a high-level planner and a low-level controller. The planner decomposes the overall goal into a series of sub-goals and decides which one the controller should address. The controller then focuses on achieving the current sub-goal by employing a specific skill. Specifically, for a certain $t$, the policy can be represented as:
@@
where $\pi_{1}$ is the planner’s policy and $\pi_{2}$ is the controller’s policy. The planner periodically (not at every step) checks whether the current sub-goal has been finished and determines whether to change policy, so $g_{t}$ is relatively static and does not change frequently.
Identifying implicit skills in offline trajectories
A critical challenge in training hierarchical agents, as described above, is the identification of skills within the offline trajectory data. To be more specific, given $\tau=(o_{i},a_{i})_{i=1}^{T}$, we want to split it into $\tau_{1}=(o_{i},a_{i})_{i=1}^{t_{1}},\tau_{2}=(o_{i},a_{i})_{i=t_{1}}^{t_{2}},\ldots,\tau_{k}=(o_{i},a_{i})_{i=t_{k-1}}^{T}$ where each sub-trajectories $\tau_{j}$ shows an independent skill $\pi(g_{j})$. Since some controllers use self-supervised methods to learn sub-goals (Lifshitz et al., 2023; Cai et al., 2023b), we can ignore the language labels of goal $g_{j}$, but must identify $t_{j}$ (i.e., the moments at which the agent adopts a different skill). For clarity, in the subsequent sections, we denote the skill adopted at a certain time step $t$ as $\pi_{t}$, corresponding to $\pi(g_{t})$.
@@ 3 Method
We first introduce our overall pipeline to build policies and long-horizon hierarchical agents based on unsegmented videos in Section 3.1. Then we introduce how we split the unsegmented demonstration into our policy and agent learning sequences in Section 3.2. We finally introduce the implementation details on Section 3.3.
@@ 3.1 Pipeline
Given a long sequence of trajectories $D={(o_{i},a_{i})}_{i=1}^{T}$ with interleaved visual observations $o_{i}$ and actions $a_{i}$, we first train an unconditioned policy $\pi_{\text{uncondition}}$ using imitation learning to predict the actions as follows:
$\min_{\theta}\sum_{t\in[1\dots T]}-\log\pi_{\text{uncondition}}(a_{t}|o_{1},\dots,o_{t}).$ (2)
When training on action label-free YouTube videos, the action labels are generated from Inverse Dynamics Model (Baker et al., 2022). We then evaluate this unconditioned policy $\pi_{\text{uncondtion}}$ on the entire dataset $D$ of long sequences. We developed an automated method called SBD to segment the complete video into a series of short segments $D_{seg}={(o_{i},a_{i})}_{i=m}^{n}$ based on these evaluation results, where $m$ and $n$ is the selected segmentation timestamps. These short segments $D_{seg}$ are then used for skill learning to obtain conditioned policies (Cai et al., 2023b; Lifshitz et al., 2023). Finally, we integrate these skills with vision language models to build hierarchical agents, enabling them to learn and follow long-horizon tasks in long video sequences (Wang et al., 2025, 2024b). We show the overall pipeline is illuminated in Figure 1.
3.2 Skill Boundary Detection
@@
[…]
@@
[…]
@@
SBD takes a long unsegmented trajectory as input and learns an unconditional model to predict the next action based on previous observations. The model does not know about implicit skill transitions. To simulate the model’s memory, a sliding window is used. At each time step $t$, the model predicts the next action and compares it with the ground truth to compute the loss. A skill transition is considered likely if the loss exceeds the average loss by a hyperparameter, GAP, or if an external indicator is true. In such cases, $t$ is marked as a boundary. The model’s memory is then cleared, and the algorithm proceeds to analyze the next sub-trajectory. The two core components of the algorithm are loss and external information.
@@ 3.2.1 Boundary with Entropy Loss
In this section, we explain the core idea of our method: why loss can indicate skill boundaries. To support this, we introduce three key assumptions about skills: skill consistency, skill confidence, and action deviance at skill transition.
The first assumption is that the skill being used does not change frequently. The parameter $K$ is a large number that determines how rare these changes are. The idea is that the agent should consistently stick to a skill unless there is a strong reason to switch to another, which would result in an increase in predictive loss for detecting a skill transition.
@@
The second assumption states that, at any given time, the agent is confident in the action it takes. This means the probability of the agent choosing an action with high confidence is very high. The parameter $c$ sets the minimum confidence level, while $\delta$ is a small value representing the frequency with which the policy might act with lower confidence. This assumption ensures that the policy reliably makes decisions based on its learned skills, rather than being uncertain in its actions. This helps the unconditional model make accurate predictions when the agent does not change its skill.
@@
Intuitively, if the agent changes its skill, the predictive probability will be low. This is because, in LABEL:eq:transition, in the first term the probability of $a_{t+1}$ under the original skill $\pi_{t}(a_{t+1}|o_{1:t+1})$ is low, and in the second term the probability of skill transition is low. On the contrary, if the agent does not change its skill, both terms will be high so the predictive probability will be high. Therefore, we have the following theorem regarding the bounds of relative predictive probability under scenarios of skill transition and non-transition.
@@
See Appendix A for the detailed proof. ∎
If $c>m$ and $(K-4)c^{2}>2$, then the lower bound of relative predictive probability under skill transition in Theorem 3.4 is greater than the upper bound of it under skill non-transition. Therefore, the theorem demonstrates that the model predicts actions less accurately when the agent changes its skill. Our action-prediction model uses the negative log-likelihood of actions $-\log P(a_{t}|o_{1:t})$ as loss. Consequently, at line 6 of Algorithm 1, when the loss exceeds the average loss by a predefined threshold, it is likely that a skill transition has occurred.
3.2.2 Boundary with Auxiliary Information
In Section 3.2.1, we demonstrate that our algorithm can identify the first distinguishable skill boundary using only observations and actions. However, some datasets include additional external information, such as privileged in-game data or VLM detections. Incorporating these auxiliary signals improves the detection of skill boundaries that are otherwise difficult to recognize through loss-based boundary detection alone. For example, in Minecraft, crafting an item does not involve abrupt changes in mouse movement, making it hard to detect boundaries purely from predictive action loss. In such cases, in-game logs of successful crafting events provide valuable hints. It is important to note that this is an optional component of our algorithm; as we will discuss in Section 4.3, it also performs well on datasets without external auxiliary information.
3.3 Implemetation Details
Unconditioned Policy and Segmentation Details
For the unconditioned policy $\pi_{\text{uncondition}}$ in Stage I, we use the pre-trained vpt-3x models (Baker et al., 2022). This is a foundational Transformer-XL (Dai et al., 2019) based model pretrained on large-scale YouTube data and finetuned on the early-game dataset using behavioral cloning. The hyperparameter GAP in Stage II is set to 18 considering the average trajectory length and semantics. As for the external information, we use event-based information such as mine_block:oak_log or use_item:torch available in the contractor dataset (Baker et al., 2022). To identify potential skill transitions, only the final event in a series of identical consecutive events is marked as positive, as it signifies a possible skill transition. For instance, if the agent chops a tree repeatedly and then crafts a table, the video should be split at the moment the agent mines the last block of wood. Further details are provided in Section B.3.
Method [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Average

### v2

# Open-World Skill Discovery from Unsegmented Demonstration Videos

@@ Abstract
Learning skills in open-world environments is essential for developing agents capable of handling a variety of tasks by combining basic skills. Online demonstration videos are typically long but unsegmented, making them difficult to segment and label with skill identifiers. Unlike existing methods that rely on random splitting or human labeling, we have developed a self-supervised learning-based approach to segment these long videos into a series of semantic-aware and skill-consistent segments. Drawing inspiration from human cognitive event segmentation theory, we introduce Skill Boundary Detection (SBD), an annotation-free temporal video segmentation algorithm. SBD detects skill boundaries in a video by leveraging prediction errors from a pretrained unconditional action-prediction model. This approach is based on the assumption that a significant increase in prediction error indicates a shift in the skill being executed. We evaluated our method in Minecraft, a rich open-world simulator with extensive gameplay videos available online. The SBD-generated segments yielded relative performance improvements of 63.7% and 52.1% for conditioned policies on short-term atomic tasks, and 11.3% and 20.8% for their corresponding hierarchical agents on long-horizon tasks, compared to random segmented baselines. Our method can leverage the diverse YouTube videos to train instruction-following agents. The project page is at https://craftjarvis.github.io/SkillDiscovery/.
footnotetext: Preprint version of the paper accepted at ICCV 2025.
@@ 1 Introduction
Random splitting [Uncaptioned image] rule-based
@@
Table 1: Comparisons between existing segmentation methods and our method SBD. Existing methods usually rely on human-designed rules, while our method is learning-based. Random splitting can result in a single skill spanning different segments or multiple skills located within one segment. Reward-driven methods require additional reward information, which is challenging for human annotators to label. Top-down methods often result in limited skill diversity and high computation costs. Bottom-up methods struggle in visually partially observable environments. [Uncaptioned image] are visual observations and [Uncaptioned image] are actions.
Most existing LLM-based instruction-following agents adopt a two-tier architecture of a planner and a controller. The planner decomposes high-level instructions into atomic skills, which the controller then maps to actions based on current observations for interaction with the environment [40, 38, 39, 26, 36, 9]. Training such agents typically involves segmenting long trajectories into atomic skill sequences, followed by separate training of the planner and controller.
@@
Learning skills from long-sequence videos is critical for building such hierarchical agents. However, real-world video data are often long-horizon, unsegmented, and with no fine-grained skill labels. In addition, the concept of a “skill” is ill-defined and varies widely across domains such as video gaming [6, 36, 19, 46], robotic control [53], and autonomous driving [37]. This ambiguity, along with the immense diversity of skills in open worlds, makes skill learning particularly challenging, especially in partially observable settings. To enable open-world skill learning from unsegmented demonstration videos, the first challenge is to segment them into semantically meaningful, self-contained skills.
@@
We list the existing segmentation methods in Tab.˜1. The naive random splitting methods [6, 27] divide videos into segments of predefined lengths (e.g., fixed length, uniform distribution). However, they do not ensure that each segment contains a distinct skill. Additionally, predefined lengths may not match the actual distribution of skill length in real world (see Sec.˜4.4). Reward-driven methods [34] discover skills through the environment’s reward signal. They cannot capture skills with no associated rewards and may split one skill into multiple segments when rewards are repeatedly gained during execution. Top-down methods [32, 4] rely on predefined skill sets from human experts. They use manual labeling or supervised learning to segment videos. Although this method can produce reasonable results, it is expensive and limited by the narrow range of predefined skills. Bottom-up methods [52, 29] use algorithms such as agglomerative clustering or byte-pair encoding [16] to split action sequences. However, they struggle in visually partially observable settings where both observations and actions must be considered. All the methods above rely on human-designed rules to segment the videos, which highlights the need for a learning-based method that can adaptively segment skills from unsegmented videos in open world.
@@
Inspired by event segmentation theory (EST) [44], which proposes that humans naturally split continuous experiences into discrete events when prediction errors in perceptual expectations rise, we introduce Skill Boundary Detection (SBD)—a method for automatically identifying potential skill boundaries in long trajectories. SBD employs a predictive model trained on a dataset of unsegmented videos to predict future actions based on past observations, effectively capturing temporal dependencies [3]. We then use this pretrained model to make predictions on unsegmented videos and compare them to the ground-truth actions. A significant increase in prediction error indicates a shift in the skill being executed (Theorem˜3.4), enabling us to detect boundaries between different skills. SBD relies on self-supervised learning, removing the need for additional human labeling. This allows SBD to utilize a wide range of YouTube videos to train instruction-following agents.
@@
We evaluate SBD in Minecraft [18, 15, 28], a rich open world simulator. First, we apply SBD to create a segmented Minecraft video dataset. We then re-train a video-conditioned policy [6] and a language-conditioned policy [27] on this dataset. We evaluate them on a diverse Minecraft skills benchmark [28]. The improved versions of these policies result in relative performance increases of 63.7% and 52.1% compared to the original ones, respectively. We also test their corresponding hierarchical agents that use behavioral cloning and in-context learning, achieving relative performance increases of 11.3% and 20.8%, respectively. These findings show the effectiveness of our proposed method for skill discovery from unsegmented demonstrations and its potential to advance open-world skill learning.
@@
Figure 1: Pipeline of our method SBD for discovering skills from unsegmented demonstration videos. Stage I: An unconditional Transformer-XL based policy model [11, 3] is pretrained on an unsegmented dataset to predict future actions (labeled by an inverse dynamics model) based on past observations using behavioral cloning. Stage II: The pretrained unconditional policy will produce a high predicted action loss when encountering uncertain observations (e.g., deciding whether to kill a new sheep) in open worlds. These timesteps should be marked as skill boundaries, indicating the need for additional instructions to control behaviors. We segment the long unsegmented videos into a series of short atomic skill demonstrations. Stage III: We train a conditional Transformer-XL based policy model on the segmented dataset to master a variety of atomic skills. Stage IV: Finally, we use hierarchical methods (a combination of vision-language models and the conditional policy) to model the long demonstration videos and follow long-horizon instructions.
@@ 2 Problem Formulation
Behavioral cloning learns a policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ from expert demonstrations, mapping observations to actions based on pairs $(o_{i},a_{i})_{i=1}^{T}$. In partially observable settings, the policy typically takes the observation history $o_{1:t}$ as input. To enable goal-directed behavior, recent works [22, 13, 41] learn goal-conditioned policies $\pi(o_{1:t},g)$, where $g$ denotes a target goal. Such goal-conditioned policies are referred to as skills, as they encode reusable behaviors for achieving specific goals.
@@
Directly learning policy $\pi(o_{1:t},g)$ is challenging in complex open-world environments like Minecraft. Goals such as build a house or mine diamond blocks involve numerous intermediate steps, making it computationally and conceptually difficult to model an agent’s behavior without breaking the goal into manageable parts. To tackle this, hierarchical agent architectures are often used [36, 12, 7, 20]. These architectures consist of a high-level planner and a low-level controller. The planner decomposes the overall goal into a series of sub-goals and decides which one the controller should address. The controller then focuses on achieving the current sub-goal by employing a specific skill. Specifically, for a certain $t$, the policy can be represented as:
@@
where $\pi_{1}$ is the planner’s policy and $\pi_{2}$ is the controller’s policy. The planner periodically (not at every step) checks whether the current sub-goal has been finished and decides whether to change it, so $g_{t}$ does not change frequently.
Identifying Implicit Skills in Trajectories
A critical challenge in training hierarchical agents, as described above, is the identification of skills within the trajectory data. To be more specific, given $\tau=(o_{i},a_{i})_{i=1}^{T}$, we want to split it into $\tau_{1}=(o_{i},a_{i})_{i=1}^{t_{1}},\tau_{2}=(o_{i},a_{i})_{i=t_{1}}^{t_{2}},\ldots,\tau_{k}=(o_{i},a_{i})_{i=t_{k-1}}^{T}$ where each sub-trajectories $\tau_{j}$ shows an independent skill $\pi(g_{j})$. Since some controllers use self-supervised methods to learn sub-goals [27, 6], we don’t have to identify $g_{j}$, but must identify $t_{j}$ (i.e., the moments at which the agent adopts a different skill). For clarity, in the subsequent sections, we denote the skill adopted at a certain time step $t$ as $\pi_{t}$, corresponding to $\pi(g_{t})$.
@@ 3 Method
We first introduce our overall pipeline to build policies and long-horizon hierarchical agents based on unsegmented videos in Section 3.1. Then we introduce how we split the unsegmented videos into skill sequences in Section 3.2. We finally introduce the implementation details in Section 3.3.
@@ 3.1 Pipeline
Given a long sequence of trajectories $D={(o_{i},a_{i})}_{i=1}^{T}$, we first train an unconditional policy $\pi_{\text{unconditional}}$ using behavioral cloning to predict the actions purely based on past observations as follows:
$\min_{\theta}\sum_{t\in[1\dots T]}-\log\pi_{\text{unconditional}}(a_{t}|o_{1:t}).$ (2)
When training on action label-free videos, the action labels are generated from an Inverse Dynamics Model [3] . We then evaluate this unconditional policy $\pi_{\text{uncondtional}}$ on the entire dataset $D$ of long sequences. Our proposed method SBD segments the complete video into a series of short segments $D_{seg}={(o_{i},a_{i})}_{i=m}^{n}$ based on the evaluation results, where $m$ and $n$ is the selected segmentation timestamps. $D_{seg}$ is then used for skill learning to obtain conditional policies [6, 27] that take additional videos or text as instructions. Finally, we integrate these skills with vision language models to build hierarchical agents, enabling them to complete long-horizon tasks [40, 39]. See the overall pipeline in Fig.˜1.
3.2 Skill Boundary Detection (SBD)
@@
2:Initialize: $\text{begin}\leftarrow 1$, $\text{loss\_history}\leftarrow[]$, $\text{boundaries}\leftarrow[]$
@@
5:  $\text{loss\_history.append}(\text{loss})$
6:  if $\text{loss}-\text{mean}(\text{loss\_history})>\text{GAP}$ or $e_{t}$ is true then
7:   $\text{boundaries.append}(t)$
@@
SBD takes a long unsegmented trajectory and the unconditional action-prediction model as input. A sliding window is used to simulate the model’s memory of past observations. At each time step $t$, we use the model to predict the next action and compare it with the ground truth to compute the loss. A skill change is considered likely if the loss exceeds the average loss by a hyperparameter, GAP, or if an external indicator is true. In such cases, $t$ is marked as a boundary. The model’s memory is then cleared, and the algorithm proceeds to analyze the next sub-trajectory.
We then explain the two core components of the algorithm: loss and external information.
@@ 3.2.1 Boundary with Entropy Loss
In this section, we explain the core idea of our method: why loss can indicate skill boundaries. Intuitively, without goal information, similar observations in an open world can inherently lead to different actions—even with perfect modeling. For example, upon seeing a sheep, both ignore and attack may be plausible depending on the (unknown) goal. Skill transitions are moments where such ambiguity is highest. Since continuing a skill is more likely than changing it, the policy learns to favor the former, making skill transitions statistically surprising. According to Bayes’ rule, when the loss of the policy increases largely, it is likely that a skill transition has occurred.
To support this, we introduce three key assumptions about skills: skill consistency, skill confidence, and action deviance at skill transition.
The first assumption is that the skill being used does not change frequently (less than $1/K$). The idea is that the agent should consistently stick to a skill unless there is a strong reason to change to another, which would result in an increase in predictive loss.
@@
The second assumption states that, at any given time, the agent is confident in the action it takes. In other words, the probability of the agent choosing an action with confidence $c$ is greater than $1-\delta$. This assumption ensures that the policy reliably makes decisions based on its learned skills, rather than being uncertain in its actions. This helps the unconditional model make accurate predictions when the agent does not change its skill.
@@
Intuitively, if the agent changes its skill, the predictive probability will be low. This is because, in LABEL:eq:transition, in the first term the probability of $a_{t+1}$ under the original skill $\pi_{t}(a_{t+1}|o_{1:t+1})$ is low, and in the second term the probability of skill transition is low. On the contrary, if the agent does not change its skill, both terms will be high, so the predictive probability will be high. Therefore, we have the following theorem regarding the bounds of relative predictive probability under scenarios of skill transition and non-transition.
@@
See Appendix˜B for the detailed proof. ∎
If $c>m$ and $(K-4)c^{2}>2$, then the lower bound of relative predictive probability under skill transition in Theorem˜3.4 is greater than the upper bound of it under skill non-transition. Therefore, the theorem demonstrates that the model predicts actions less accurately when the agent changes its skill. Our action-prediction model uses the negative log-likelihood of actions $\log P(a_{t}|o_{1:t})$ as loss. Consequently, at line 6 of Algorithm˜1, when the loss exceeds the average loss by a predefined threshold, it is likely that a skill transition has occurred.
3.2.2 Boundary with External Information
In Sec.˜3.2.1, we demonstrate that SBD can identify the first distinguishable skill boundary using only observations and actions. However, some datasets include additional external information such as privileged in-game data or VLM annotations [guo2025seed1, comanici2025gemini]. Incorporating these auxiliary signals improves the detection of skill boundaries that are otherwise difficult to recognize through loss-based detection alone. For example, in Minecraft, crafting an item does not involve abrupt changes in mouse movement, making it hard to detect boundaries purely with predictive action loss. In such cases, in-game logs of successful crafting events provide valuable hints. It is important to note that this is an optional component of SBD; as we will discuss in Sec.˜4.3, it also performs well on datasets without external auxiliary information.
Task
use
furnace
hunt
sheep
sleep
in bed
use
torch
use
boat
use
bow
collect
stone
collect
seagrass
collect
wood
find and
collect wood
collect
dirt
collect
grass
Average